OpenStudy (anonymous):

What are two negative fractions with has a product of 5/8?

OpenStudy (aaliyahfairgood):

-5/8 and -1/8

OpenStudy (anonymous):

So that would be -5/8 * -1/8 = 5/8?

OpenStudy (aaliyahfairgood):

yes!

OpenStudy (anonymous):

Okay, thanks!

OpenStudy (aaliyahfairgood):

no problem. a negative times a negative is always a positive answer

OpenStudy (anonymous):

I also need help with this: What is the error and what is the correction? -1/4 / 3/2 = -4/1 x 3/2 = -12/2 = -6

OpenStudy (aaliyahfairgood):

is it -1/4 times 3/2

OpenStudy (aaliyahfairgood):

oh I see

OpenStudy (anonymous):

No it's -1/4 divided by 3/2

OpenStudy (aaliyahfairgood):

-1/4 divided by 3/2 flip 3/2 to get 2/3, then multiply, also known as multiplying by the reciprocal -1/4 * 2/3

OpenStudy (aaliyahfairgood):

-1/4 x 2/3 = 2/12 = 1/6

OpenStudy (aaliyahfairgood):

did that help

OpenStudy (anonymous):

Yes! thank you for your help

OpenStudy (aaliyahfairgood):

your very welcome

OpenStudy (triciaal):

the first problem is incorrect

OpenStudy (triciaal):

it is true that a negative times a negative is a positive but with fractions you multiply the denominator 8*8 = 64

OpenStudy (triciaal):